# Toward Efficient and Robust Behavior Models for Multi-Agent Driving Simulation

### v2

@@ I INTRODUCTION
Given the high cost and potential risks of deploying automated vehicles in real-world environments, simulation is indispensable for the research and development of advanced driver assistant systems. Key applications include generating diverse scenarios that may be difficult or dangerous to recreate in the real world [suo2023mixsim, chitta2024scenario_generation2, amini2020RAL_RL_edge_case], predicting human driving behavior [zhang2023trafficbots, bhattacharyya2022modeling, Konstantinidis2024IV], and training or evaluating behavior models [apple2025self-play, cornelisse2025self-play].
A key requirement for multi-agent driving simulation is exhibiting realistic behavior to reduce the sim-to-real gap. For that, behavior models map observations to actions. They are typically trained via Imitation Learning (IL), e. g., [bansal2018chauffeurnet, lu2023imitation], which learns directly from real data but yields less robust policies, or via Reinforcement Learning (RL), e. g., [apple2025self-play, chen2019model, konstantinidis2023ITSC, arief2024RAL_IS_meta_rl], which improves robustness through exploration in simulation but depends on a reward function, which is unknown for real-world drivers.
Instead, we employ Adversarial Inverse Reinforcement Learning (AIRL) [fu2018airl] to reconstruct a reward signal from real traffic data and simultaneously learn a behavior model that maximizes this reward signal using standard RL. Since the reconstructed reward signal can differ across experiments, it may unintentionally affect the robustness of the model. To address this, we introduce an adaptive reward transformation that balances robustness and realism during training.
@@
The second requirement is achieving high execution speed to enable efficient scaling. While prior works, e. g., [gulino2023waymax, kazemkhani2024gpudrive, scibior2021torchdrivesim, apple2025self-play], have focused on accelerating the simulation backbone, our focus is on optimizing the executed behavior model. Commonly, agent-centric observations (see Figure 1a) are used, e. g., [Konstantinidis2024IV, kazemkhani2024gpudrive, apple2025self-play, cornelisse2025self-play], where nearby instances, such as agents and static map elements, are represented from each agent’s perspective. However, generating and encoding these local observations for every agent at each simulation step imposes a significant computational burden. Alternatively, scene-centric observations (see Figure 1b) represent all instances in a global coordinate frame, e. g., [suo2021trafficsim, ngiam2021scene]. While this offers efficient encoding, it comes at the potential sacrifice of pose-invariance, requiring the model to implicitly learn variations in agent positions and orientations.
@@ II Related Work
Realistic Behavior Simulation: Early approaches to simulating vehicle behavior include log-replay [bergamini2021simnet, gulino2023waymax], where trajectories recorded in the real world are simply replayed, rule-based models [kesting2007general, kesting2010enhanced, gulino2023waymax], as well as Behavior Cloning (BC) [lu2023imitation, bansal2018chauffeurnet], an open-loop IL method, where a mapping from observations to actions is learned from a dataset using supervised learning. Due to its simplicity and the availability of large-scale datasets, the latter remains widely used. Each approach, however, has its limitations: Replayed trajectories are non-reactive, thus prohibiting interactions with a simulated ego vehicle. Rule-based models often fail to capture the complexity of real-world behaviors. BC performs poorly in rare or unseen situations (e. g., near-collisions or very aggressive merges). Even worse, when deploying such a model in a closed-loop simulation, small deviations from expert driving accumulate, leading the model into unfamiliar states - known as covariate shift [spencer2021feedback].
@@
For learning more robust behavior models, RL is commonly employed [chen2019model, cornelisse2025self-play, apple2025self-play, konstantinidis2023ITSC, arief2024RAL_IS_meta_rl]. Since realistically simulating traffic situations is inherently a multi-agent problem, self-play RL has emerged as an effective solution. In [konstantinidis2023ITSC], all agents are controlled by a single shared policy, which is trained jointly through self-play. To enable applying the model to a diverse set of traffic situations, a flexible, graph-based input representation with vectorized map features is adopted. Similarly, in [apple2025self-play] and [cornelisse2025self-play], self-play RL has been applied on a large scale, demonstrating that robust and naturalistic behavior emerges from it. The resulting behavior model generalizes well, even in out-of-distribution scenarios. Notably, in [apple2025self-play], the behavior model is trained in randomly initialized traffic situations, driving more than $1.6$ billion km during training. The resulting behavior model is zero-shot evaluated on three autonomous driving benchmarks, achieving state-of-the-art performance without ever seeing human data during training.
@@
However, achieving realistic behavior with RL requires a carefully tuned reward function that precisely guides the agents toward human-like driving. Recent works [Konstantinidis2024IV, sun2024modelling, weaver2024RAL_residualAIL, bhattacharyya2022modeling] address this issue by reconstructing a reward signal from real-world data. In [Konstantinidis2024IV], AIRL is used, where a discriminator is trained to distinguish real from simulated behavior, assigning higher scores to more realistic samples. As the goal is to drive as realistically as possible, the output of the discriminator is then used as a reward signal for RL training. Similarly, we train our model using AIRL, but introduce an adaptive reward transformation to balance robustness and realism.
@@
Enhanced inference speed: Many existing works on multi-agent driving simulation focus on optimizing the simulation backbone, i. e., the engine responsible for executing actions and producing next-step observations. This is commonly done by implementing the simulator in PyTorch [scibior2021torchdrivesim, apple2025self-play], JAX [gulino2023waymax], or C++ using CUDA [kazemkhani2024gpudrive] to support in-graph compilation for hardware (GPU/TPU) acceleration. However, we argue that often, not the simulation engine, but the executed behavior model is the limiting factor. We therefore focus on optimizing the executed behavior model for multi-agent simulation.
@@
Although coordinated scene-centric agent behavior has already been investigated for single-shot motion forecasting, e. g., [ngiam2021scene, su2022narrowing], to the best of our knowledge, TrafficSim [suo2021trafficsim] is the only work on learning a scene-centric multi-agent behavior model for closed-loop simulation. In [suo2021trafficsim], a global map is rasterized and encoded using a CNN. Then, local map features are extracted via Rotated Region of Interest Align and fused with the agent features. Lastly, a joint decoder model, realized as a message passing network, processes all agent features jointly, enabling coordination between them. This scene-centric representation effectively leads to more efficient simulations, as the static environment needs to be encoded only once, allowing the reuse of its encoded tokens in subsequent simulation steps. However, due to the global coordinate frame, this comes with the potential sacrifice of pose-invariance, often leading to performance degradation when the global coordinate frame changes.
To address this, recent single-shot motion prediction methods, e. g., [zhang2024simpl, zhou2023query, zhang2023hptr, shi2024mtr++], adopt instance-centric representations, encoding each agent and map element in its own local frame. These pose-invariant features can be reused across all target agents, with relative positions enabling symmetric encoding from any agent’s perspective.
@@ III Method
For learning and evaluating behavior models, a simulation framework is required. In our in-house simulation framework, each vehicle is assigned a route to follow. Similarly, existing works use waypoints [gulino2023waymax, suo2023mixsim] or goal points [kazemkhani2024gpudrive, apple2025self-play]. Vehicles terminate upon colliding or leaving the road.
@@ III-A Reinforcement Learning
In our multi-agent setting, we apply a shared policy across all agents, trained via self-play RL. Specifically, we use Generalized Advantage Estimation (GAE) [Schulmann15] for estimating $A(o_{k},a_{k})$ and Proximal Policy Optimization (PPO) [Schulmann17] for optimizing the policy. We refer to [apple2025self-play] for a detailed description of how self-play RL is applied to the task of learning behavior models for simulating driver behavior.
@@ III-B Adversarial Inverse Reinforcement Learning
RL requires defining a reward function that accurately captures the incentive structure of real-world driving. Instead of manually defining such a reward function, we use AIRL [fu2018airl] to reconstruct a surrogate reward signal from real data, given as $\mathcal{D}=\{(o_{1},a_{1}),(o_{2},a_{2}),\dots\}$. In AIRL, an additional discriminator model $D_{\phi}$ is trained to distinguish generated from real samples, outputting the probability $D_{\phi}(o,a)\in\left[0,1\right]$ for the observation-action pair being real, i. e., stemming from $\mathcal{D}$. The policy is trained via RL using the surrogate reward
@@
which guides the policy to taking realistic actions. Following [Konstantinidis2024IV], to smooth the decision boundary during discriminator training, we add random noise to the actions in $\mathcal{D}$, matching the standard deviation of the policy.
@@
Adaptive Reward Transformation: Typically, the discriminator classifies samples correctly, leading to a negative expected value for $\tilde{r}(o,a)$. This often drives agents to terminate as fast as possible to avoid constantly accumulating penalties. To mitigate this, in [Konstantinidis2024IV] a positive constant $c=5$ is added to $\tilde{r}$, shifting its expected value into the positive range. Intuitively, this adjustment encourages survival without altering the policy’s optimal behavior w. r. t. the discriminator.
@@ III-C Instance-Centric Observations
Rather than adopting agent-centric observations, as prior works do, e. g., [bhattacharyya2022modeling, apple2025self-play, cornelisse2025self-play], we employ instance-centric observations, representing each instance, i. e., the $N_{a}$ agents and $N_{p}$ map elements, in its respective local coordinate frame.
For static map elements, such as pedestrian crossings or lane boundaries, we use a vectorized representation [vectornet]. Each map element is approximated by a polyline with a maximum total length of $10\text{\,}\mathrm{m}$. A polyline is composed of multiple vectors, with each being described by a set of features $\mathbf{v}=\left[\mathbf{v}_{\mathrm{start}},\mathbf{v}_{\mathrm{end}},\mathbf{v}_{\mathrm{type}}\right]^{\intercal}$, including the vector’s start and end point, as well as a one-hot encoding for the polyline type. For polylines, the origin of the instance-centric coordinate frame is given as the mean of all polyline points, with the x-axis being aligned to the tangential direction of the polyline.
@@
Intuitively, the local coordinate frames serve as anchor poses for the individual instances, with $\mathbf{p}_{i}$ being the origin and $\alpha_{i}$ being the heading of the $i$-th instance relative to an arbitrary global coordinate frame. Following [zhang2024simpl, shi2024mtr++], to capture the spatial relationship between two instances $i$ and $j$, we define a coordinate transformation from the local frame of $i$ to that of $j$. Specifically, the relative pose of instance $j$ in the frame of instance $i$ is represented as
@@ III-D Behavior Model Details
Instance Feature Encoders: We begin by processing the observation into a set of latent instance-tokens. Following the approach introduced in [vectornet], the $L$ vectors of a polyline are encoded through multiple layers of message-passing:
@@
For that, we define the relative positional encoding between a target agent $i\in\{0,\cdots,N_{a}\}$ and any of the combined instance tokens $Z=\left[Z^{\mathrm{(agent)}},\;Z^{\mathrm{(map)}}\right]$, indexed by $j$, as $\mathbf{c}_{i\rightarrow j}=\mathbf{r}_{i\rightarrow j}\oplus\mathds{1}^{(\mathrm{agent})}_{j}\oplus\mathds{1}^{(\mathrm{route})}_{j}$, where $\oplus$ denotes concatenation. The binary indicators $\mathds{1}^{(\mathrm{agent})}_{j}$ and $\mathds{1}^{(\mathrm{route})}_{j}$ specify whether instance $j$ is an agent, or a map polyline that is part of the route that agent $i$ should follow, respectively. These indicators are appended here rather than to the vector features $\mathbf{v}$, as each agent may follow a different route, making a global encoding at the vector level infeasible. Finally, following the approach proposed in [perez2018film], we compute pairwise encodings by applying a feature-wise linear transformation:
@@
Similarly to [chen2024drivingwithllms, zhang2024simpl], we then refine the instance tokens of simulated target agents through multiple Perceiver layers [jaegle2021perceiver]. Each layer consists of a Multi-Head Cross Attention (MHCA) module with a skip connection, followed by layer normalization, an MLP, also with a skip connection, and another layer normalization. The MHCA is expressed as:
@@ IV Experiments
Datasets: We evaluate our approach on two datasets: 1) the publicly available INTERACTION dataset [zhan2019interaction], containing 38 25538\,255 trajectories ($4.15\text{\,}\%$ VRUs) across $11$ locations; and 2) an in-house dataset commercially licensed by DeepScenario, comprising 88 54088\,540 trajectories ($43.59\text{\,}\%$ VRUs) across $8$ locations. Both datasets were captured using drones, ensuring high-quality data, and include highly interactive scenarios such as merging lanes, unsignalized intersections, and roundabouts. We use $20\text{\,}\%$ of the recordings per location for validation and $30\text{\,}\%$ for testing. A more compact version of the DeepScenario dataset [dhaouadi2025deepscenario] is publicly available and can be used to obtain similar results.
@@
    LateFusionMLP [cornelisse2025self-play]: Following [apple2025self-play, cornelisse2025self-play, kazemkhani2024gpudrive], this compact agent-centric model consists solely of MLP s and max-pooling operations. We adopt the public implementation [cornelisse2025self-play], replacing its discrete action decoder with ours to support continuous actions and training it within our framework for realistic behavior modeling.
@@
    GraphAIRL [Konstantinidis2024IV]: A more sophisticated agent-centric model that leverages a vectorized scene representation [vectornet] and attention-based interaction modeling. We evaluate two variants: 1) trained with $c=5$, as proposed in [Konstantinidis2024IV], and 2) trained with our proposed adaptive reward offset, defined in (3).
@@
Training Details: All models are trained using the AdamW optimizer [loshchilov2017adamw] with a batch size of $1024$. Learning rates are set to $0.0002$ for the policy model and $0.0001$ for the discriminator, both decayed by a factor of $10$ during the final $30\text{\,}\%$ of training. For RL, we use advantage normalization. The models are trained for 10 00010\,000 epochs with a discount factor $\gamma=0.95$, GAE parameter $\lambda=0.95$, and target rewards $\tilde{r}_{\mathrm{target}}=19$ and $\tilde{r}_{\mathrm{target}}=11$ for INTERACTION and DeepScenario, respectively. For the policy model, the observation radius is set to $50\text{\,}\mathrm{m}$. Since the number of observed instances grows quadratically with the observation range, we reduce the discriminator’s observation radius to $30\text{\,}\mathrm{m}$, allowing it to focus primarily on the target agent and its direct surroundings.

### v3

@@ I INTRODUCTION
Given the high cost and potential risks of deploying automated vehicles in real-world environments, simulation is indispensable for the research and development of advanced driver assistant systems. Key applications include generating diverse scenarios that may be difficult or dangerous to recreate in the real world [33, 8, 1], predicting human driving behavior [37, 5, 20], and training or evaluating behavior models [10, 9].
A key requirement for multi-agent driving simulation is exhibiting realistic behavior to reduce the sim-to-real gap. For that, behavior models map observations to actions. They are typically trained via Imitation Learning (IL), e. g., [3, 22], which learns directly from real data but yields less robust policies, or via Reinforcement Learning (RL), e. g., [10, 6, 19, 2], which improves robustness through exploration in simulation but depends on a reward function, which is unknown for real-world drivers.
Instead, we employ Adversarial Inverse Reinforcement Learning (AIRL) [12] to reconstruct a reward signal from real traffic data and simultaneously learn a behavior model that maximizes this reward signal using standard RL. Since the reconstructed reward signal can differ across experiments, it may unintentionally affect the robustness of the model. To address this, we introduce an adaptive reward transformation that balances robustness and realism during training.
@@
The second requirement is achieving high execution speed to enable efficient scaling. While prior works, e. g., [14, 16, 27, 10], have focused on accelerating the simulation backbone, our focus is on optimizing the executed behavior model. Commonly, agent-centric observations (see Figure 1a) are used, e. g., [20, 16, 10, 9], where nearby instances, such as agents and static map elements, are represented from each agent’s perspective. However, generating and encoding these local observations for every agent at each simulation step imposes a significant computational burden. Alternatively, scene-centric observations (see Figure 1b) represent all instances in a global coordinate frame, e. g., [32, 23]. While this offers efficient encoding, it comes at the potential sacrifice of pose-invariance, requiring the model to implicitly learn variations in agent positions and orientations.
@@ II Related Work
Realistic Behavior Simulation: Early approaches to simulating vehicle behavior include log-replay [4, 14], where trajectories recorded in the real world are simply replayed, rule-based models [17, 18, 14], as well as Behavior Cloning (BC) [22, 3], an open-loop IL method, where a mapping from observations to actions is learned from a dataset using supervised learning. Due to its simplicity and the availability of large-scale datasets, the latter remains widely used. Each approach, however, has its limitations: Replayed trajectories are non-reactive, thus prohibiting interactions with a simulated ego vehicle. Rule-based models often fail to capture the complexity of real-world behaviors. BC performs poorly in rare or unseen situations (e. g., near-collisions or very aggressive merges). Even worse, when deploying such a model in a closed-loop simulation, small deviations from expert driving accumulate, leading the model into unfamiliar states - known as covariate shift [29].
@@
For learning more robust behavior models, RL is commonly employed [6, 9, 10, 19, 2]. Since realistically simulating traffic situations is inherently a multi-agent problem, self-play RL has emerged as an effective solution. In [19], all agents are controlled by a single shared policy, which is trained jointly through self-play. To enable applying the model to a diverse set of traffic situations, a flexible, graph-based input representation with vectorized map features is adopted. Similarly, in [10] and [9], self-play RL has been applied on a large scale, demonstrating that robust and naturalistic behavior emerges from it. The resulting behavior model generalizes well, even in out-of-distribution scenarios. Notably, in [10], the behavior model is trained in randomly initialized traffic situations, driving more than $1.6$ billion km during training. The resulting behavior model is zero-shot evaluated on three autonomous driving benchmarks, achieving state-of-the-art performance without ever seeing human data during training.
@@
However, achieving realistic behavior with RL requires a carefully tuned reward function that precisely guides the agents toward human-like driving. Recent works [20, 31, 34, 5] address this issue by reconstructing a reward signal from real-world data. In [20], AIRL is used, where a discriminator is trained to distinguish real from simulated behavior, assigning higher scores to more realistic samples. As the goal is to drive as realistically as possible, the output of the discriminator is then used as a reward signal for RL training. Similarly, we train our model using AIRL, but introduce an adaptive reward transformation to balance robustness and realism.
@@
Enhanced inference speed: Many existing works on multi-agent driving simulation focus on optimizing the simulation backbone, i. e., the engine responsible for executing actions and producing next-step observations. This is commonly done by implementing the simulator in PyTorch [27, 10], JAX [14], or C++ using CUDA [16] to support in-graph compilation for hardware (GPU/TPU) acceleration. However, we argue that often, not the simulation engine, but the executed behavior model is the limiting factor. We therefore focus on optimizing the executed behavior model for multi-agent simulation.
@@
Although coordinated scene-centric agent behavior has already been investigated for single-shot motion forecasting, e. g., [23, 30], to the best of our knowledge, TrafficSim [32] is the only work on learning a scene-centric multi-agent behavior model for closed-loop simulation. In [32], a global map is rasterized and encoded using a CNN. Then, local map features are extracted via Rotated Region of Interest Align and fused with the agent features. Lastly, a joint decoder model, realized as a message passing network, processes all agent features jointly, enabling coordination between them. This scene-centric representation effectively leads to more efficient simulations, as the static environment needs to be encoded only once, allowing the reuse of its encoded tokens in subsequent simulation steps. However, due to the global coordinate frame, this comes with the potential sacrifice of pose-invariance, often leading to performance degradation when the global coordinate frame changes.
To address this, recent single-shot motion prediction methods, e. g., [36, 39, 38, 28], adopt instance-centric representations, encoding each agent and map element in its own local frame. These pose-invariant features can be reused across all target agents, with relative positions enabling symmetric encoding from any agent’s perspective.
@@ III Method
For learning and evaluating behavior models, a simulation framework is required. In our in-house simulation framework, each vehicle is assigned a route to follow. Similarly, existing works use waypoints [14, 33] or goal points [16, 10]. Vehicles terminate upon colliding or leaving the road.
@@ III-A Reinforcement Learning
In our multi-agent setting, we apply a shared policy across all agents, trained via self-play RL. Specifically, we use Generalized Advantage Estimation (GAE) [25] for estimating $A(o_{k},a_{k})$ and Proximal Policy Optimization (PPO) [26] for optimizing the policy. We refer to [10] for a detailed description of how self-play RL is applied to the task of learning behavior models for simulating driver behavior.
@@ III-B Adversarial Inverse Reinforcement Learning
RL requires defining a reward function that accurately captures the incentive structure of real-world driving. Instead of manually defining such a reward function, we use AIRL [12] to reconstruct a surrogate reward signal from real data, given as $\mathcal{D}=\{(o_{1},a_{1}),(o_{2},a_{2}),\dots\}$. In AIRL, an additional discriminator model $D_{\phi}$ is trained to distinguish generated from real samples, outputting the probability $D_{\phi}(o,a)\in\left[0,1\right]$ for the observation-action pair being real, i. e., stemming from $\mathcal{D}$. The policy is trained via RL using the surrogate reward
@@
which guides the policy to taking realistic actions. Following [20], to smooth the decision boundary during discriminator training, we add random noise to the actions in $\mathcal{D}$, matching the standard deviation of the policy.
@@
Adaptive Reward Transformation: Typically, the discriminator classifies samples correctly, leading to a negative expected value for $\tilde{r}(o,a)$. This often drives agents to terminate as fast as possible to avoid constantly accumulating penalties. To mitigate this, in [20] a positive constant $c=5$ is added to $\tilde{r}$, shifting its expected value into the positive range. Intuitively, this adjustment encourages survival without altering the policy’s optimal behavior w. r. t. the discriminator.
@@ III-C Instance-Centric Observations
Rather than adopting agent-centric observations, as prior works do, e. g., [5, 10, 9], we employ instance-centric observations, representing each instance, i. e., the $N_{a}$ agents and $N_{p}$ map elements, in its respective local coordinate frame.
For static map elements, such as pedestrian crossings or lane boundaries, we use a vectorized representation [13]. Each map element is approximated by a polyline with a maximum total length of $10\text{\,}\mathrm{m}$. A polyline is composed of multiple vectors, with each being described by a set of features $\mathbf{v}=\left[\mathbf{v}_{\mathrm{start}},\mathbf{v}_{\mathrm{end}},\mathbf{v}_{\mathrm{type}}\right]^{\intercal}$, including the vector’s start and end point, as well as a one-hot encoding for the polyline type. For polylines, the origin of the instance-centric coordinate frame is given as the mean of all polyline points, with the x-axis being aligned to the tangential direction of the polyline.
@@
Intuitively, the local coordinate frames serve as anchor poses for the individual instances, with $\mathbf{p}_{i}$ being the origin and $\alpha_{i}$ being the heading of the $i$-th instance relative to an arbitrary global coordinate frame. Following [36, 28], to capture the spatial relationship between two instances $i$ and $j$, we define a coordinate transformation from the local frame of $i$ to that of $j$. Specifically, the relative pose of instance $j$ in the frame of instance $i$ is represented as
@@ III-D Behavior Model Details
Instance Feature Encoders: We begin by processing the observation into a set of latent instance-tokens. Following the approach introduced in [13], the $L$ vectors of a polyline are encoded through multiple layers of message-passing:
@@
For that, we define the relative positional encoding between a target agent $i\in\{0,\cdots,N_{a}\}$ and any of the combined instance tokens $Z=\left[Z^{\mathrm{(agent)}},\;Z^{\mathrm{(map)}}\right]$, indexed by $j$, as $\mathbf{c}_{i\rightarrow j}=\mathbf{r}_{i\rightarrow j}\oplus\mathds{1}^{(\mathrm{agent})}_{j}\oplus\mathds{1}^{(\mathrm{route})}_{j}$, where $\oplus$ denotes concatenation. The binary indicators $\mathds{1}^{(\mathrm{agent})}_{j}$ and $\mathds{1}^{(\mathrm{route})}_{j}$ specify whether instance $j$ is an agent, or a map polyline that is part of the route that agent $i$ should follow, respectively. These indicators are appended here rather than to the vector features $\mathbf{v}$, as each agent may follow a different route, making a global encoding at the vector level infeasible. Finally, following the approach proposed in [24], we compute pairwise encodings by applying a feature-wise linear transformation:
@@
Similarly to [7, 36], we then refine the instance tokens of simulated target agents through multiple Perceiver layers [15]. Each layer consists of a Multi-Head Cross Attention (MHCA) module with a skip connection, followed by layer normalization, an MLP, also with a skip connection, and another layer normalization. The MHCA is expressed as:
@@ IV Experiments
Datasets: We evaluate our approach on two datasets: 1) the publicly available INTERACTION dataset [35], containing 38 25538\,255 trajectories ($4.15\text{\,}\%$ VRUs) across $11$ locations; and 2) an in-house dataset commercially licensed by DeepScenario, comprising 88 54088\,540 trajectories ($43.59\text{\,}\%$ VRUs) across $8$ locations. Both datasets were captured using drones, ensuring high-quality data, and include highly interactive scenarios such as merging lanes, unsignalized intersections, and roundabouts. We use $20\text{\,}\%$ of the recordings per location for validation and $30\text{\,}\%$ for testing. A more compact version of the DeepScenario dataset [11] is publicly available and can be used to obtain similar results.
@@
    LateFusionMLP [9]: Following [10, 9, 16], this compact agent-centric model consists solely of MLP s and max-pooling operations. We adopt the public implementation [9], replacing its discrete action decoder with ours to support continuous actions and training it within our framework for realistic behavior modeling.
@@
    GraphAIRL [20]: A more sophisticated agent-centric model that leverages a vectorized scene representation [13] and attention-based interaction modeling. We evaluate two variants: 1) trained with $c=5$, as proposed in [20], and 2) trained with our proposed adaptive reward offset, defined in (3).
@@
Training Details: All models are trained using the AdamW optimizer [21] with a batch size of $1024$. Learning rates are set to $0.0002$ for the policy model and $0.0001$ for the discriminator, both decayed by a factor of $10$ during the final $30\text{\,}\%$ of training. For RL, we use advantage normalization. The models are trained for 10 00010\,000 epochs with a discount factor $\gamma=0.95$, GAE parameter $\lambda=0.95$, and target rewards $\tilde{r}_{\mathrm{target}}=19$ and $\tilde{r}_{\mathrm{target}}=11$ for INTERACTION and DeepScenario, respectively. For the policy model, the observation radius is set to $50\text{\,}\mathrm{m}$. Since the number of observed instances grows quadratically with the observation range, we reduce the discriminator’s observation radius to $30\text{\,}\mathrm{m}$, allowing it to focus primarily on the target agent and its direct surroundings.